Analysis and Interpretation of Profitability
Balance sheets and income statements for Target Corporation follow.

Income Statement
For Fiscal Years Ended ($ millions) 2006 2005 2004
Sales $ 51,271 $ 45,682 $ 40,928
Credit card revenues 1,349 1,157 1,097
Total revenues 52,620 46,839 42,025
Cost of sales 34,927 31,445 28,389
Selling, general and administrative expenses 11,185 9,797 8,657
Credit card expenses 776 737 722
Depreciation and amortization 1,409 1,259 1,098
Earnings before interest and income taxes 4,323 3,601 3,159
Net interest expense 463 570 556
Earnings before income taxes 3,860 3,031 2,603
Provisions for income taxes 1,452 1,146 984
Net earnings $ 2,408 $ 1,885 $ 1,619
Balance Sheet
($ millions, except footnotes) January 28, 2006 January 29, 2005
Assets
Cash and cash equivalents $ 1,648 $ 2,245
Credit card receivables 5,666 5,069
Inventory 5,838 5,384
Other current assets 1,253 1,224
Total current assets 14,405 13,922
Property and equipment
Land 4,449 3,804
Buildings and improvements 14,174 12,518
Fixtures and equipment 3,219 2,990
Computer hardware and software 2,214 1,998
Construction-in-progress 1,158 962
Accumulated depreciation (6,176) (5,412)
Property and equipment, net 19,038 16,860
Other noncurrent assets 1,552 1,511
Total assets $ 34,995 $ 32,293
Liabilities and shareholders' investment
Accounts payable $ 6,268 $ 5,779
Accrued and other current liabilities 2,567 1,937
Current portion of long-term debt and notes payable 753 504
Total current liabilities 9,588 8,220
Long-term debt 9,119 9,034
Deferred income taxes 851 973
Other noncurrent liabilities 1,232 1,037
Shareholders' investment
Common stock 73 74
Additional paid-in-capital 2,121 1,810
Retained earnings 12,013 11,148
Accumulated other comprehensive income (loss) (2) (3)
Total shareholders' investment 14,205 13,029
Total liabilities and shareholders' equity $ 34,995 $ 32,293


HINT: For Sales use "Total revenues" for your computations, when applicable.

(a) Compute net operating profit after tax (NOPAT) for 2006. Assume that the combined federal and statutory rate is: 38.3%. (Round your answer to the nearest whole number.)
2006 NOPAT = ? $million

(b) Compute net operating assets (NOA) for 2006 and 2005.
2006 NOA = ? $million
2005 NOA = ? million

(c) Compute Target’s RNOA, net operating profit margin (NOPM) and net operating asset turnover (NOAT) for 2006. (Do not round until final answer. Round two decimal places. Do not use NOPM x NOAT to calculate RNOA.)
2006 RNOA = ? %
2006 NOPM = ? %
2006 NOAT = ?

(d) Compute net nonoperating obligations (NNO) for 2006 and 2005.
2006 NNO = $Answer million
2005 NNO = $Answer million

(e) Compute return on equity (ROE) for 2006. (Do not round until final answer. Round answer two decimal places.)
2006 ROE = Answer%

(f) Infer the nonoperating return component of ROE for 2006. (Use answers from above to calculate. Round your answer to two decimal places.)
2006 nonoperating return = Answer%

(g) Which of the following statements reflects the best inference we can draw from the difference between Target's ROE and RNOA?

ROE>RNOA implies that Target's equity has grown faster than its NOA.

ROE>RNOA implies that Target has taken on too much financial leverage.

ROE>RNOA implies that Target is able to borrow money to fund operating assets that yield a return greater than its cost of debt; the excess accrues to the benefit of Target's stockholders.

ROE>RNOA implies that Target has increased its financial leverage during the period.
Answer #1

NOPAT = operating profit x (1-tax rate)

Or in other words, NOPAT = EBIT(1-tax rate)

For 2006, EBIT = $4323

Therefore for 2006 NOPAT = 4323x(1-0.383)

= $2667.29

2 . Net operating assets are those assets of a Company which are directly related to its operations and business less liabilities directly related to its business.

Thus, we can say that Net operating asset = total assets - total liabilities - financial asset + financial liabilities

For 2005 & 2006 calculation of net operating asset will be as follows :-

Particulars 2006 2005
Total assets 34995 32293
Less : total liabilities (20790) (19264)
Less : construction in progress (1158) (962)
Add : current portion of long term debt and notes payable 753 504
Add : long term debt 9119 9034
Net operating asset $22919 $21605

C . Return on net operating asset = NOPAT/average NOA

= 2667.29/{(22919+21605)/2}x100 = 11.98%.

NOPM = = NOPAT/sales x100 = 2667.29/51271x100 = 5.20%.

NOAT = revenue / average NOA

= 51271/{(22919+21605)/2} = 2.30

D . Net non operating obligation = excess of non operating debt over investment

For 2005 & 2006

Particulars 2006 2005
Current portion of long term debt and notes payable 753 504
Add : long term debt 9119 9034
Less : construction in progress (1158) (962)
Net non operating obligation $8714 $8576

E . Return on equity = net income / average shareholder investment

= 2408/{(14205+13029)/2}x100 = 2408/13617x100

= 17.68%.

F. Non operating Return component of ROE = ROE - RNOA

= 17.68-5.20 = 12.48%.

F. Note thst ROE>RNOA implied that target is able to borrow money to fund operating asset that yields a return greater than it's it's cost of debt and as a result the excess accrues to the benefit of target's stockholders . Thus , correct answer is (c)